Can you help me solve this problem?
The perimeter is 38 Meters and its area is 84 square meters, find its dimension.
Hope you can help me :)

Answer by Cromlix(4381)

Hi there,
Perimeter = 2 x Length + 2 x Width
Area = Length x Width
...........
Area = Length x Width
84 = Length x Width
Width = 84/Length
Substitute W = 84/L into perimeter
Perimeter = 2 x Length + 2 x Width
38 = 2L + 2(84/L)
38 = 2L + 168/L
Multiply thro'out by L
38L = 2L^2 + 168
Sort into equation
(ax^2 + bx + c = 0)
2L^2 - 38L + 168 = 0
Divide through by 2
L^2 - 19L + 84 = 0
Factorise
(L - 7)(L - 12) = 0
Therefore L = 7m W = 12m
or L = 12m W = 7m
Hope this helps:-)